# Unleashing the Potential of Large Language Models for Text-to-Image Generation through Autoregressive Representation Alignment

### v2

@@ 4.2.1 Visual encoder selection on alignment
4.2.2 Alignment mechanism
We compare two strategies for integrating visual representations: (1) aligning features to a fixed <REP> token at the start of the generated sequence, and (2) aligning to the hidden state of <HYBNEXT>, a hybrid token interleaved at every generation step. As shown in Fig. 2, <HYBNEXT> yields superior performance. We argue that <HYBNEXT> allows for comprehensive traversal of every token during training sampling, ensuring that each token is effectively constrained by external global representations. In contrast, <REP> suffers from the “attention sink” [27, 57] effect, where attention to the fixed token decays over long sequences, leading to degraded outputs. This leads to the following key insight: Takeaway 2. Aligning visual representations to a hybrid token interleaved at each generation step <HYBNEXT> is more effective than using a fixed token <REP>, as it prevents attention decay and ensures consistent constraint by external representations.
CLS Avgpool FID $\downarrow$ MS-SSIM $\uparrow$ CLIP-Score $\uparrow$
5.30 0.405 0.4532
6.56 0.387 0.4434
5.93 0.385 0.4465
Figure 3: Visualization of different methods and ARRA in the MIMIC-CXR dataset for the chest X-ray generation task.
4.2.3 Alignment depth
These results lead to a key insight: Takeaway 3. Performing representation alignment at earlier LLM layers (e.g., layer 1) provides sufficient global semantic guidance while allowing deeper layers to focus on capturing high-frequency details.
Table 5: FID, IS, MS-SSIM comparisons of different methods on the MIMIC-CXR, DeepEyeNet, and ImageNet datasets, respectively.
Chest-Xray: MIMIC-CXR Fundus: DeepEyeNet Natural: ImageNet
Method FID $\downarrow$ IS $\uparrow$ MS-SSIM $\uparrow$ CLIP-Score $\uparrow$ FID $\downarrow$ MS-SSIM $\uparrow$ CLIP-Score $\uparrow$ Method Epochs FID $\downarrow$
Original SD V2-1[45] 71.68 2.365 0.128 0.2333 166.45 0.141 0.2164 LlamaGen- [^1] M 100 9.328
DreamBooth SD [46] 60.40 2.269 0.270 0.3693 80.70 0.352 0.3061 LlamaGen- [^1] M 300 7.283
MINIM [54] 15.62 2.468 0.317 0.4423 59.71 0.333 0.2862 LlamaGen-343M 100 5.192
UniXGen [22] 30.75 2.437 0.361 0.4128 - - - LlamaGen-343M 300 4.294
LLM-CXR [23] 5.88 2.134 0.395 0.4374 - - - ARRA-Base- [^1] M 100 8.710
Chameleon [50] 7.11 2.498 0.383 0.4460 38.37 0.341 0.3234 ARRA-Base- [^1] M 300 6.653
ARRA 5.30 2.587 0.405 0.4532 35.01 0.376 0.3354 ARRA-Base-343M 100 4.982
ARRA-Adapt 4.15 2.746 0.422 0.4576 34.70 0.392 0.3389 ARRA-Base-343M 300 3.971
Figure 4: ARRA-Base improves generation of LlamaGen.
Figure 5: Convergence trends of the baseline and ARRA models on the MIMIC-CXR (left) and ImageNet (right) datasets.
4.2.4 Feature aggregation strategy
To investigate how different types of features extracted from the visual encoder affect the generation performance, we investigate three strategies for aggregating features from a single visual encoder: (1) the [CLS] token representation, (2) average pooling of all image patch representations, and (3) a concatenation of both representations.
As shown in Table 4, the [CLS] token representation yields optimal performance. We attribute this superiority to the [CLS] token’s ability to aggregate global visual information through self-attention mechanisms, providing a compact yet comprehensive representation for cross-modal alignment. This finding aligns with [40], which establishes that [CLS] tokens capture global representations while patch tokens focus on local features. This distinction further suggests that autoregressive image generation models particularly benefit from external global semantic representations as guidance.
These results lead to a key insight: Takeaway 4. The [CLS] token representation in foundation models effectively aggregates global visual information, providing comprehensive guidance for cross-modal alignment.
4.3 Main Comparison Results
Based on our component analysis, we finalize our ARRA framework with BioMedCLIP and CLIP for visual alignment on medical datasets and natural dataset, respectively. We use [CLS], [HYBNEXT], early-layer (layer 1) as our alignment strategy. We now benchmark our models against state-of-the-art diffusion models (Stable Diffusion [45], DreamBooth [46], MINIM [54]) and autoregressive models (LLM-CXR [23], UniXGen [22], Chameleon [50]). More implementation details are placed in Appendix 2.
Quantitative and qualitative performance. We evaluate generation quality using FID (visual fidelity), MS-SSIM (structural similarity), and CLIP-Score (text-image alignment). As shown in Fig. 3 and Table 5, our model achieves outstanding performance in both visual quality and semantic alignment aspects. Among the baseline methods, LLM-CXR [23] achieves respectable FID and MS-SSIM scores but exhibits poor semantic alignment. Conversely, MINIM [54] achieves excellent CLIP-Score performance but suffers from significant image distortion. These conclusions are also reflected in the visual results shown in Fig. 3. In contrast, both our ARRA model and the ARRA-Adapt model show excellent performance in visual fidelity and semantic alignment, achieving the best results. Specifically, compared to the LLM-CXR model, which achieves the best FID among the baselines, the ARRA model reduces the FID from 5.88 to 5.30 and significantly improves the CLIP-Score from 0.4374 to 0.4532. The ARRA-Adapt model further reduces the FID to 4.15 and increases the CLIP-Score to 0.4576. In comparison to the Chameleon model in the baseline, which serves as foundation for our ARRA model, ARRA achieves a significant reduction in FID by 25.5%.
Additionally, our models ensure excellent alignment in fine-grained details, such as lesion location and severity. For example, when generating images for first report, Chameleon model incorrectly positions “right pleural effusion”, and it fails to adequately represent the severity of “mild enlargement”. In comparison, our ARRA and ARRA-Adapt model guarantees this fine-grained understanding and achieves high-quality generation. More visualizations results are in Appendix 5.
Accelerate convergence. In Fig. 5, we present the convergence trends of the baseline model and the model enhanced with ARRA framework. The experimental results show that applying ARRA at different layers results in faster convergence compared to baseline, particularly in the early stages of training. This indicates that the global constraints introduced by ARRA’s alignment mechanism effectively guide the model along the correct learning trajectory, thereby accelerating convergence. Additional visual comparisons of convergence trends are provided in Appendix 5.
Generalization Ability. We validate our framework on another popular autoregressive generative model, LlamaGen [48], and define models that use our framework as ARRA-Base model. We employ the FID as the metrics and evaluate the performances of our models on ImageNet dataset. As shown in Table 5 and Fig. 4, ARRA-Based achieved better generation performance than the baseline in both  [^1] M and 343M models. Fig. 5 indicates that our method beat LlamaGen model during the entire training process. Through experimental results, we show that the ARRA framework can be applied to various autoregressive models and is suitable for different datasets in multiple domains. More ablation study can be found in Appendix 4. More implementation details and visualizations are shown in Appendix 5.
5 Conclusion
We propose Autoregressive Representation Alignment (ARRA), a framework that enhances autoregressive image generation by injecting external visual representations during training. This approach enriches global semantic understanding while maintaining the model’s original autoregressive paradigm during generation. Experiments on medical and natural image generation tasks demonstrate ARRA’s versatility, offering a cost-effective framework for training autoregressive text-to-image generation models. Our work bridges the gap between multimodal domains and provides novel insights into modeling unified multimodal generation.

### v4

@@ 5 More Ablation study
Figure 2: Visualization of alignment with features extracted from different encoders.
Figure 3: Visualization of different lesion localizations and lesion levels.
Figure 4: More visual comparison of images generated by the baseline and the ARRA model (with alignment) during the iteration process.
6 More visualizations on MIMIC-CXR
We present additional visual results from experiments on MIMIC-CXR, including:
  • Visual comparisons of ARRA with other methods on X-ray image generation tasks (Fig. 1);
  • Visualizations of baseline models using different pretrained encoders for alignment (Fig. 2);
  • Visualization of different lesion localizations and lesion levels. (Fig. 3);
  • Convergence performance visualizations of baseline models and those using the ARRA framework (Fig. 4).
7 More implementation and results in LAION-COCO and ImageNet
7.1 Datasets
We collect 2.4M high-quality image-text pairs from the LAION-COCO (schuhmann2022laion) dataset as our training data. We additionally collect 10K samples as a test set. All images are center-cropped to 256×256. We use subset provied by GoT (fang2025got), filtering with similarity >> 0.3. For ImageNet (imagenet), we conduct experiments on the ImageNet 256×256 conditional generation benchmark.
7.2 Implementation details
ARRA-Base. We apply the ARRA framework on LlamaGen (sun2024autoregressivemodelbeatsdiffusion) to test the generalization ability of our model. In our experiments, we use a batch size of 32 with a learning rate of 1e-4 and a vocabulary size of 16,384. For the visual foundation model, we use CLIP-L (radford2021CLIP), where the extracted feature, our target representation, has the shape (batch_size, 512). In the  [^1] M model, the dimension of the hidden states is 768, while in the 343M model, the dimension of the hidden states is 1024. We use an MLP layer as the feature summarization tool to reduce the dimension of hidden states, making it possible to be aligned with our target. We set $\lambda=1$ and use cosine similarity to calculate the loss between the target representation and the hidden states.
Figure 5: More visualizations of ARRA-Base-3.1B on the LAION-COCO dataset.
Table 6: ARRA-Base in image generation task across the ImageNet Dataset (classifier-free guidance=2.00)
Method Para epochs data FID $\downarrow$ IS $\uparrow$
LlamaGen  [^1] M 100 0.26M 9.328 126.4
LlamaGen  [^1] M 150 0.26M 8.421 142.7
LlamaGen  [^1] M 200 0.26M 8.010 145.4
LlamaGen  [^1] M 250 0.26M 7.580 161.8
LlamaGen  [^1] M 300 0.26M 7.283 156.7
LlamaGen  [^1] M 300 1.28M 5.464 193.6
LlamaGen 343M 100 0.26M 5.192 187.8
LlamaGen 343M 150 0.26M 4.666 206.7
LlamaGen 343M 200 0.26M 4.419 202.0
LlamaGen 343M 300 0.26M 4.294 218.6
LlamaGen 343M 300 1.28M 4.327 286.6
ARRA-Base  [^1] M 100 0.26M 8.710 133.0
ARRA-Base  [^1] M 150 0.26M 7.873 142.2
ARRA-Base  [^1] M 200 0.26M 7.483 145.8
ARRA-Base  [^1] M 250 0.26M 6.737 162.7
ARRA-Base  [^1] M 300 0.26M 6.653 159.0
ARRA-Base 343M 100 0.26M 4.982 192.1
ARRA-Base 343M 150 0.26M 4.544 200.8
ARRA-Base 343M 200 0.26M 4.182 206.8
ARRA-Base 343M 300 0.26M 3.971 219.1
7.3 More experiments results
This is a more comprehensive experiment results of ARRA-Base on the Imagenet datasets (Table 6), where we prove that ARRA enhances the generation ability of LlamaGen. ARRA works for both  [^1] M and 343M models, indicating that this method is scalable.
7.4 Visualizations
In this section, we present additional visualizations (Fig. 5, 6, 7, 8, 9) on the LAION-COCO and ImageNet datasets, comparing ARRA-Base and LlamaGen.
Figure 6: More visualizations of ARRA-Base-343M on ImageNet dataset.
Figure 7: More visualizations of LlamaGen-343M on ImageNet dataset.
Figure 8: More visualizations of ARRA-Base- [^1] M on ImageNet dataset.
Figure 9: More visualizations of LlamaGen- [^1] M on ImageNet dataset.